OpenStudy (anonymous):

Solve the equation: 26 + 8y - 18 = 0

OpenStudy (anonymous):

where you are stuck???

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

do u know how to start?

OpenStudy (anonymous):

I'm not very good with this kind of math

OpenStudy (anonymous):

this is simplest math ever

OpenStudy (anonymous):

just tell yes or no?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok what is 26-18?

OpenStudy (anonymous):

subtract the constant first

OpenStudy (anonymous):

I know I have to minus or plus somthing :p

OpenStudy (anonymous):

what you got??

OpenStudy (anonymous):

8

OpenStudy (anonymous):

ok so u have 8y-8 =0 left understand?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

now 8y +8 = 0 can you find y now???

OpenStudy (anonymous):

yeah sry 8y +8=0

OpenStudy (anonymous):

do u undestand how we got this?

OpenStudy (anonymous):

8y divided by somthing

OpenStudy (anonymous):

ok so now subtract by 8 on both sides of the equation...can u do that?

OpenStudy (anonymous):

take the constant 8 to right side what you get ??

OpenStudy (anonymous):

??? minus 8 from 8 =0?

OpenStudy (anonymous):

yes...and 0-8=?

OpenStudy (anonymous):

0y+0=0?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hahaha its 8y + 8 - 8 = 0- 8

OpenStudy (anonymous):

nooo!! 8y +8 -8 =0 -8 so 8y=-8 so y=-1 get it?